OpenStudy (anonymous):

integral (2n^2 -1)/(3n^5 +2n+1) dn from 1 to infinity

OpenStudy (amistre64):

thats a rather odd looking rational expression

OpenStudy (amistre64):

it mimics 2/3n^3, but that prolly doesnt help

OpenStudy (amistre64):

by parts looks doable; if not lengthy ....

OpenStudy (amistre64):

cant even see that

OpenStudy (anonymous):

i got 1/3 but the answer on wolfram is 0.206

OpenStudy (amistre64):

what steps did you take? numerical perhaps?

OpenStudy (mr.math):

I assume you're using this integral to test if the series \(\sum_{n=1}^n \frac{2n^2 -1}{3n^5 +2n+1}\) converges, right?

OpenStudy (mr.math):

from n=1 to infinity*

OpenStudy (amistre64):

if so then the comparison test might be good: \[\lim_{n->inf}\ \frac{2n^2 -1}{3n^5 +2n+1}*\frac{3n^3}{2}=1\] ....well thats inconclusive

OpenStudy (amistre64):

\[\frac{2n^2 -1}{3n^5 +2n+1}\ ?\le?\frac{2}{3n^3}\] \[6n^5 -3n^2\ ?\le?6n^5 +4n+2\] \[0\ \le\ 7n+2;\ n\ge0\]its true so if 2/3n^3 converges; the lesser function has to as well

OpenStudy (amistre64):

the 2/3 is rather moot; so 1/n^3 is a p-series, where the (p)ower is > 1 and so it converges